OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? A. 4x - y = 13 B. 4x + y = 13 C. 4x + y = -13 I got A ?

OpenStudy (anonymous):

close.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok, let take steps 4x +y-2=0--> 4x+y=2 ok?

OpenStudy (anonymous):

now, replace the point (4,-3) to the left hand side you have 4*4 +(-3) = 13, so that 4x \(\color{red}{+} \) y =13

OpenStudy (anonymous):

Oh so messed up on a step. it's supposed to be the plus sign instead of the minus?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Well thank you! I understand how to do it now. @Bebong